Find the sum of the infinite series ∑_(n=1)^(+∞)▒1/(k^2+5k+6)


1
Expert's answer
2021-05-12T03:04:27-0400

Given series is-


"\sum_{k=1}^{\infty}\dfrac{1}{k^2+5k+6}"


"=\sum_{k=1}^{\infty} \dfrac{1}{(k+3)(k+2)}



 \\[9pt]=\sum_{k=1}^{\infty} (\dfrac{1}{(k+2)}-\dfrac{1}{(k+3)}



\\[9pt]=(\dfrac{1}{3}-\dfrac{1}{4})+\dfrac{1}{4}-\dfrac{1}{5})+\dfrac{1}{5}-\dfrac{1}{6}+.....



\\[9pt] =(\dfrac{1}{3}+\dfrac{1}{4}+\dfrac{1}{5}+.....)-(\dfrac{1}{4}+\dfrac{1}{5}+.....)



\\[9pt]=\dfrac{1}{3}"